7. Consider the two functions: f(x, a)-40 - 3xa 9(2) 10 + 7x (a) Find the x value of the point where the two equa- (b) Find the value of the functions at the point where erwya tions intersect (in terms of the variable a) the two equations intersect (in terms of the variable a) (c) Take the partial derivate of f with respect to , and with respec t to a. d) What are the values of these derivatives when r - 3 2, which can be written as and a (3.2) and 3, 2 (e) Next, calculate these two numbers v1 = (3.2) yatt er v2 = Con (3.2) T(3,2)(f) Finally write out these equations in terms of a and r and simplify, erwyt Эт f (x, a) a) Эт За f (x, a)
Answer #1

Answer 7

a)

f(x,a) = 40 - 3xa

g(x) = 10 + 7x

when these functions intersects f(x,a) = g(x)

=> 40 - 3xa = 10 + 7x

=> 30 = x(3a + 7)

=> x = 30/(3a + 7)

b) When they intersect.

f(a) = g(a) = (30a + 280)/(3a + 7)

c)

Here

f(x,a) = 40 - 3xa

=>

3 DT

Also

af 0 3 Oa

d)

As calculated above

rac{partial f}{partial x} = rac{partial f}{partial a} = -3

for any (x,a)

Hence when x = 3 and and a = 2

rac{partial f(3,2)}{partial x} = rac{partial f(3,2)}{partial a} = -3